Three very long current carrying wires are placed at the corners of an equilateral triangle as shown below. Wire 1 and wire 2 are fixed at their locations well above the ground using some anchors. Wire 3 with linear mass density 40 g/m is being levitated magnetically above the ground due to the other two wires. (a) Find the direction of current in the wire 3 to achieve this. (b) Find the magnitude of the current in wire 3 (I3) to make it levitate above the ground if I1 = I2 = 75 A.

4.0 cm x Х I 12 13
Answer #1

SOLUTION 4.0cm 0 Iz 5 , Equilateral triangle . 0 = 60° F32 Iz Part (a) For given condition, current in wire 3 must be such th4x 10² kg/m F32 = Fonce per unit length on wire 3 due to wire 1. llo I2 I3 zad [4x107] [75] [ Iz 2[ 4x162] =[37.5x105] 53 Nmi for given condition: Fnet w (64.952x10-5) 13 = 39-2x16 2 or, I3 = 39.2x 10-2 64. gS2x105 - 603.5226 A